Fermi National Laboratory


Accelerator Update for October 19, 2005 - October 21, 2005

Wednesday October 19
The day shift began with the complex recovering from the power glitch, with Tevatron (TeV) experts conducting studies, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store, with NuMI off for chiller work, and with the MiniBooNE and the Meson MCenter experiments also waiting for beam.

At 10:20 AM, Operations reported that while they were trying to circulate beam in the Main Injector (MI), the P150 transfer line suffered a vacuum burst. Operators started closing vacuum valves to try and contain the burst, but some valves wouldn't close. Eventually the vacuum recovered.

At 11:01 AM, water technicians reported that they found a LCW leak while they were filling the HPTF LCW system in the Meson Detector building. (The High Powered Test Facility is located in the Meson East beam line area.)

Operations began preparing the Recycler at 10:24 AM to stash antiprotons. The stash ended. Pbar began stacking at 11:04 AM.

Operations established beam to the Meson MCenter experiment at 11:25 AM.

A Switchyard (SY) expert began a study at 12:39 PM.

At 1:48 PM, techs began work on a MI-50 regulator chassis. This work shut off MI beam.

The MI work ended and Pbar resumed stacking at 2:47 PM.

The Recycler began preparing at 6:01 PM to stash.

At 6:04 PM, an expert and an operator investigated a problem with a MI RF station (MIRF1). They reported at 6:19 PM that the station was down and it appeared to be a problem in the tunnel. They bypassed the station.

At 7:01 PM, Operations reported that they had been having a problem with Booster. beam for NuMI. They called in an expert.

The Recycler stash ended and Pbar resumed stacking at 7:08 PM.

Operations put the TeV into shot setup at 7:50 PM.

Operations established Recycler only store 4460 at 9:22 PM with an initial luminosity of 84.9E30. Pbar resumed stacking.

At 10:45 PM, Operations reported that they weren't getting beam to the Debuncher. Operators continued to tune the beam through the AP1 transfer line.

Thursday October 20
The midnight shift began with Operations monitoring store 4460, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Recycler began preparing at 1:23 AM to stash.

At 2:27 AM, Operations reported that the stash was on hold due initially to problems with Pbar and MI, but then the big problem was that the stash sequencer faulted. Operations contacted a Recycler expert to help get the RF into the correct state.

At 2:50 AM, with help from an expert, Operations ended the stash, reconfigured the Recycler, and Pbar resumed stacking. MCenter also resumed taking beam.

Around 3 AM, Operations reported that beam had stopped making it into the Accumulator. Operators investigated.

At 3:40 AM, Operations reported that a Debuncher dipole bus power supply (D:IB) ground faulted.

Operators reported at 4:28 AM that though there was no alarm for D:IB until it actually tripped, the plots showed that this device was causing the Pbar problems. An expert was called in to make a tunnel access. The Recycler stashed what antiprotons there were left in Pbar.

The power supply expert reported that he had come in and turned on D:IB. At 4:47 AM, he reported that D:IB had been on for fifteen minutes and was operating normally. Operations put Pbar back into stacking mode.

Operations established beam to MCenter at 4:49 AM.

NuMI asked to be turned off at 5 AM so they could continue with their Horn chiller work.

At 5:15 AM, CDF called about beam in the abort gap. Operations called the Tevatron Electron Lens (TEL) expert who advised operators on how to manipulate the TEL's timing. CDF put their silicon detector in standby while operators adjusted the timing.

At 5:17 AM, Operations reported that D:IB didn't cycle properly. An operator and an expert investigated. Operations called in more people to help with this problem at 5:38 AM.

MiniBooNE asked to have their beam turned off at 5:52 AM for a scheduled network outage.

At 6 AM, D0 reported that they were seeing too large a halo. This was during the TEL tuning.

An operator completed the TEL timing adjustment and CDF turned their detector back on at 6:06 AM.

MiniBooNE resumed taking beam at 6:31 AM.

At 8:13 AM, Operations reported that the TeV beam halo was still high. Operations called in the TEL expert and he had the beam halo lowered to an acceptable level by 9:42 AM.

A power supply expert reported at 10:18 AM that D:IB was ready to run.

Pbar resumed stacking at 10:49 AM. A Pbar expert then resumed his BPM studies. The study ended at 11:15 AM.

NuMI reported at 4:16 PM that they had finished their chiller work for the day and would like beam.

NuMI experts began a target scan at 4:29 PM.

Operations established beam to the Meson MTest experiment at 5:07 PM.

The Recycler began preparing at 6:01 PM to stash. MTest quit using beam for the night. The stash ended and Pbar resumed stacking at 6:59 PM. MCenter resumed taking beam.

TeV experts began an end-of-store study at 11 PM.

Friday October 21
The midnight shift began with Operations monitoring store 4460 and the stash, with TeV experts conducting studies, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler began preparing at 1:56 AM to stash.

At 2:30 AM, an AP1 power supply (M:HT105) tripped off and wouldn't reset. Operators discovered that there was no spare regulator for this supply. Operations called in experts and then tuned around this supply. This problem delayed the stash.

The TeV studies ended 4:16 AM.

Operations terminated store 4460 at 4:38 AM.

The stash was completed at 4:50 AM. Pbar and MCenter resumed normal operations at 5:14 AM.

Operations put the TeV into shot setup at 6:22 AM.

At 6:33 AM, the lack of readbacks caused by a TeV sector A2 crate failure and then a sequencer problem put a halt to shot setup. Operations called in an expert. Operations resumed sending beam to Pbar and MCenter.

Operations put the TeV through a dry squeeze at 7:33 AM. A TeV expert had modified the sequencer.

The Plans for Friday and the Weekend
The plans for today and this weekend are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/21/2005   email Fermilab